### v1

# HMT-Grasp: A Hybrid Mamba-Transformer Approach for Robot Grasping in Cluttered Environments

Abstract
Robot grasping, whether handling isolated objects, cluttered items, or stacked objects, plays a critical role in industrial and service applications. However, current visual grasp detection methods based on Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) struggle to adapt across various grasping scenarios due to the imbalance between local and global feature extraction. In this paper, we propose a novel hybrid Mamba-Transformer approach to address these challenges. Our method improves robotic visual grasping by effectively capturing both global and local information through the integration of Vision Mamba and parallel convolutional-transformer blocks. This hybrid architecture significantly improves adaptability, precision, and flexibility across various robotic tasks. To ensure a fair evaluation, we conducted extensive experiments on the Cornell, Jacquard, and OCID-Grasp datasets, ranging from simple to complex scenarios. Additionally, we performed both simulated and real-world robotic experiments. The results demonstrate that our method not only surpasses state-of-the-art techniques on standard grasping datasets but also delivers strong performance in both simulation and real-world robot applications.
I Introduction
Robotic grasping has become increasingly important in industrial production and service applications [1, 2]. Tasks such as automated warehouse picking and household robots rely on accurate grasping to handle objects effectively. While early methods using hand-crafted features performed well in simple, controlled environments, modern robotic systems face more complex scenarios with diverse object shapes and unpredictable scenarios, making grasping a greater challenge [3].
Deep learning techniques, particularly Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have greatly advanced computer vision tasks such as image classification, object detection, and segmentation [4, 5]. Researchers have leveraged these techniques to improve robotic grasp detection by utilizing their ability to extract crucial visual information. Early CNN-based grasp detectors [6, 7, 8, 9, 10] demonstrated notable improvements over hand-crafted methods, adapting to a wider variety of objects. However, CNNs focus mainly on local feature extraction, often missing the global context needed for grasping in more complex scenarios. To address this issue, Vision Transformers [11, 12] have been introduced in grasping tasks, offering enhanced global feature representation by capturing information across the entire input image. For instance, Wang et al. [13] and Dong et al. [14] employed transformers to encode target scenes for robot grasping detection. While transformer-based approaches enhance global feature representation, they tend to decrease concern on local details as the model depth increases [15].
A more recent approach, Vision Mamba (ViM), balances global and local information more effectively than ViTs by incorporating Visual State Space (VSS) models [16]. While ViM improves spatial detail capture and focuses more on local features than transformers, its local feature extraction remains less refined than CNNs. In this work, we propose a hybrid Mamba-Transformer approach, integrating the strengths of Vision Mamba [17], transformers [18], and CNNs to enhance robotic grasp detection across various scenarios. By combining outputs from parallel CNNs and transformers as input to the Mamba encoder, our approach balances global and local feature representations, addressing the limitations of each model. Experimental results show that our method outperforms state-of-the-art techniques on multiple grasping datasets and delivers superior performance in both simulation and real-world robotic applications.
The main contributions of this paper can be summarized as follows:
1.
We propose a novel hybrid Mamba-Transformer approach that integrates the strengths of Vision Mamba, transformers, and CNNs to improve robotic grasp detection. To the best of our knowledge, we are the first to explore the performance of the Mamba model in robotic grasping tasks.
2.
We provide a comprehensive evaluation of our method on a diverse set of grasping datasets, demonstrating its effectiveness across various scenarios.
II Related Work
This section mainly reviews recent advancements in robotic grasp detection. Current research on robot grasp detection is primarily categorized into two approaches: CNN-based methods and ViT-based methods. Additionally, we provide a brief overview of recent progress with Vision Mamba in various domains.
II-A Grasp Detection
CNN-Based Methods: Convolutional Neural Networks (CNNs) have been widely applied in robotic visual grasping tasks [9]. Early work by Lenz et al. introduced a two-step cascaded system that employed two deep neural networks to predict the grasp pose of objects. In this approach, numerous candidate grasps are first generated and then re-evaluated in a second step, which significantly slows down the detection process [13]. Subsequent research has leveraged CNNs to enhance object grasping performance by predicting grasp rectangles. Redmon et al. proposed a single-stage regression model that enabled real-time grasp detection [8]. To further improve accuracy, a ResNet-50-based grasp detector was introduced, demonstrating superior performance on the Cornell RGB-D dataset [19]. Datasets such as Jacquard [20] and Multi-Object [21] have contributed to enhancing the accuracy and generalization of CNN models in grasp detection. In addition, the Generative Grasping CNN (GG-CNN), designed to work with depth images, improved pixel-wise grasp quality in real-time scenarios. GR-ConvNet [10], proposed by Kumra et al., used n-channel inputs to detect antipodal grasps, achieving high accuracy on the Cornell and Jacquard datasets. Despite the effectiveness of CNN-based methods in grasp detection, a significant drawback is their tendency to lose global context due to the inherent structure of stacked convolutional layers [13].
ViT-Based Methods: Motivated by the success of Vision Transformers (ViTs) in object detection and image classification, researchers are increasingly exploring their potential for predicting object grasp poses in robotic tasks [13, 14, 22, 23]. Wang et al. [13] were among the first to propose the use of Swin-Transformers for visual grasp detection, employing them as both encoders and decoders with skip connections to generate pixel-level grasp representations. Building on this, Dong et al. [14] also utilized Swin-Transformers for encoding, coupled with fully convolutional decoders to enhance grasp detection performance. In a more recent effort, Yenicesu et al. [23] introduced LITV2, a lightweight ViT model, to achieve real-time grasp detection. Compared to CNN-based methods, ViT-based approaches have demonstrated improved grasp detection accuracy. However, as the depth of transformer blocks increases, local-level information tends to gradually diminish [15].
II-B Visual Mamba
Recently, Mamba, a novel State Space Model (SSM) initially developed for natural language processing (NLP) [24], has shown strong potential for efficiently modeling long sequences. This success has sparked increasing interest in applying SSMs to visual tasks. In particular, Mamba has been successfully implemented in various computer vision applications. For example, ViM [25], a Mamba-based architecture, operates on image patch sequences using a 1D causal scanning mechanism. However, the 1D scanning approach faces challenges in fully capturing global representations. To address this limitation, VMamba [26] introduces 2D-Selective-Scan, transforming images into sequences of patches for improved global feature extraction.
Inspired by these advancements, SegMamba [27] and VM-UNet [28] have successfully integrated visual state space models into encoders for medical image segmentation, contributing to notable performance improvements in this domain. Despite these developments, no prior work has explored the application of SSMs in robotic grasp detection. Given the strengths of mamba in capturing both global information and local details, alongside the complementary capabilities of transformers and CNNs, we designed a Mamba-Transformer approach to predict robotic grasp poses. Our method demonstrates superior grasp success rates compared to state-of-the-art transformer- and CNN-based approaches.
III Method
In this section, we introduce a hybrid Mamba-Transformer architecture for robotic grasp detection. The overall structure of the model is shown in Fig. 1.
III-A Grasp Representation
We define the grasp representation , based on prior works [13, 9], as key parameters describing a successful grasp. Using a parallel-plate gripper constrained to the - plane, the grasp is represented by:
where are the center coordinates of the grasp rectangle, and are the width and height, is the rotation angle. For an image with channels and dimensions , the predicted grasp is:
here, are the coordinates, is the rotation angle, is the grasp width, and is the grasp quality. To convert image coordinates into the robot’s reference frame, we apply a series of transformations:
where transforms from image space to the camera frame using intrinsic parameters, and converts to the robot frame. The grasp information in the image is stored as a grasp map:
where represent the angle, width, and quality for each pixel. The hybrid Mamba-Transformer framework analyzes the n-channel image to determine the optimal grasping pose, focusing on predicting the grasp location, orientation, and width.
III-B Grasp Encoder
The input tensor to the architecture has dimensions , where is the batch size, is the number of channels, and represent the height and width of the image. The tensor dimensions change as it passes through different layers of the model.
III-B1 Stem-Downsample Block
Initially, the input tensor is processed by the Stem-Downsample Block, consisting of two Conv2d layers, each followed by BatchNorm (BN) and SiLU activations:
III-B2 Parallel Convolutional-Transformer Blocks
Following the StemDown Block, the output is processed through parallel Transformer and Convolutional Blocks.
Transformer Blocks: To capture global information, we employ the Swin Transformer block [11], which uses Window-based Multi-Head Self-Attention (W-MSA) and Shifted Window-based MSA (SW-MSA), followed by LayerNorm (LN) and Multi-Layer Perceptron (MLP).
Convolutional Blocks: To refine local features, a DoubleConv block is used, consisting of Conv2D, BN, and ReSiLU activations, where ReSiLU combines the outputs of ReLU and SiLU activations:
III-B3 Downsampling
To merge the information from the Transformer and Convolutional Blocks, we first compute the average of and , denoted as . This is followed by a downsampling network consisting of MaxPooling and DoubleConv layers for further refinement:
The resulting fused tensor undergoes an additional DoubleConv operation, ensuring efficient downsampling and feature extraction.
III-B4 Mamba Blocks
The fused and downsampled tensor with dimension is processed by the Mamba Block, inspired by the VM-UNet architecture [28]. The Mamba Block consists of LayerNorm (LN), Linear Projection (LP), Conv2d, SiLU activation, and 2D-Selective-Scan (SS2D). The computation is expressed as:
In this formulation, the input undergoes Layer Normalization before being processed by a series of operations, including a linear projection, convolution, and the SiLU activation. The resulting feature map is further refined using the 2D-Selective-Scan (SS2D) mechanism, followed by an element-wise combination of the two branches. A final residual connection is applied to retain the input information and improve gradient flow during training. This structure aims to effectively capture spatial details and enhance the overall feature representation for subsequent stages of the network.
III-C Grasp Decoder
After processing by the grasp encoder, the feature representation is passed to a DoubleConv block for further feature extraction. This DoubleConv block consists of Conv2D, Batch Normalization (BN), and ReSiLU activation layers, similar to the ones used in the grasp encoder. During this step, the input and output dimensions remain unchanged. The output of the DoubleConv block is combined with the output from the Mamba block. Mathematically, this can be expressed as:
The first fused output () is passed through an upsampling block. After upsampling, the output with dimensions is processed by another DoubleConv block for refinement:
The first upsampled output is then summed with the output from the parallel convolutional-transformer block in the grasp encoder, followed by an additional upsampling operation.
This process ensures the preservation and refinement of spatial details through feature refinement, skip connections, and upsampling. The resulting tensor with dimensions is further upsampled to generate pixel-level heatmap images for grasp prediction.
III-D Grasp Prediction
In the final output of the HMT-Grasp detector, pixel-level heatmap images are generated for grasp representation. These include a grasp confidence map , a gripper angle map , and a gripper width map , all matching the input image size (224x224). Grasp posture estimation is formulated as a regression task, where the model learns the mapping to minimize the difference between predicted grasp heatmaps and ground truth , with as the input data. The loss function is defined as:
where represents the number of samples, and denotes the ground truth for each sample.
The optimal grasp location is determined by selecting the position with the highest grasp confidence , as defined by:
where represents the chosen grasp location. The corresponding gripper angle and width are then extracted from the respective heatmaps.
IV Experiments
In this section, we present a comprehensive evaluation of our method through experiments on the Cornell, Jacquard, and OCID-Grasp datasets. We assess the impact of skip connections and Mamba blocks on performance and compare our approach to state-of-the-art methods. Additionally, we validate our method in both simulation and real-world robotic experiments. The grasp network has been implemented using PyTorch 2.0.1 with cuDNN 8.5.0.96 and CUDA 11.8. The model is trained end-to-end on an Nvidia RTX 4060 Ti GPU with 16GB of memory.
IV-A Dataset and Experiment setup
This section outlines the datasets used in the evaluation and provides relevant training details.
IV-A1 Cornell Dataset and Training
The Cornell dataset is one of the most widely used benchmarks for robotic grasping. It consists of 885 images of 250 real, graspable objects with a total of 8019 annotated grasps. The dataset contains RGB-D images and can be split using two different methods: image-wise and object-wise.
Given the relatively small size of the Cornell dataset, we employ 5-fold cross-validation to train and evaluate our model. To ensure robust performance, data augmentation is applied during training. The model is optimized using the Adam optimizer, with a learning rate initialized at 0.0002 and a batch size of 8. Training is carried out over 100 epochs to obtain the final model weights.
IV-A2 Jacquard Dataset and Training
The Jacquard grasp dataset is a large-scale dataset designed for robotic grasping tasks, comprising over 54,000 synthetic RGB-D images of more than 11,000 distinct objects. It includes approximately 1.1 million labeled grasps, generated using a physics-based simulation environment, providing diverse and comprehensive data for training grasping models.
We employ image-wise splitting to train and test our network, where 90% of the data is used for training, and the remaining 10% is reserved for testing. While we utilize the Adam optimizer, similar to the Cornell dataset training, we adjust the learning rate to 0.001 and set the batch size to 8. The network is trained over 200 epochs to obtain the final model weights.
IV-A3 OCID Dataset and Training
The OCID-Grasp dataset is a large-scale, real-world dataset created for robotic grasping in cluttered environments. It contains 1763 selected images with over 75,000 hand-annotated grasp candidates, classified into 31 object categories. The dataset is designed to capture challenging grasping scenarios involving a variety of objects in cluttered settings.
We utilize 5-fold cross-validation, as in the Cornell dataset training, to train the networks and report average performance across the folds. The Adam optimizer is employed with a learning rate of 0.0004 and a batch size of 8. The network is trained for 70 epochs to achieve the final model weights.
IV-B Evaluation Metrics
Similar to previous studies [9, 10, 13], this work evaluates grasp detection using the following criteria:
Angle Difference: The orientation difference between the predicted grasp and the ground-truth is less than 30.
Jaccard Index: The Jaccard index between the predicted grasp and the ground-truth is greater than 25%, calculated as:
here, and denote the predicted and ground-truth grasp rectangles, respectively, with the intersection and union represented as and .
IV-C Performance Comparison with State-of-the-Art Methods
To demonstrate the effectiveness of our approach, we compared it against state-of-the-art methods under consistent training and testing settings across the Cornell, Jacquard, and OCID-Grasp datasets.
IV-C1 Grasping Performance on the Cornell Dataset
| Method | Input | Accuracy (%) | Time (ms) | |
| IW | OW | |||
| GG-CNN [9] | D | 73.0 | 69.0 | 19 |
| GPRN [29] | RGB | 88.7 | - | 200 |
| HybridGrasp [30] | RGB-D | 93.2 | 89.1 | - |
| GR-ConvNet [10] | RGB | 96.6 | 95.5 | 19 |
| GR-ConvNet [10] | RGB-D | 97.7 | 96.6 | 20 |
| TF-Grasp [13] | D | 95.2 | 94.9 | 41.1 |
| TF-Grasp [13] | RGB | 96.78 | 95.0 | 41.3 |
| TF-Grasp [13] | RGB-D | 97.99 | 96.7 | 41.6 |
| HRG-Net [31] | D | 99.43 | 96.8 | 52.6 |
| HRG-Net [31] | RGB | 98.50 | 96.7 | 53.0 |
| HRG-Net [31] | RGB-D | 99.50 | 97.5 | 53.7 |
| AE-GDN [32] | RGB | 97.2 | 96.4 | 24.6 |
| AE-GDN [32] | RGB-D | 98.9 | 97.9 | 26.4 |
| DCSFC-Grasp [33] | RGB-D | 99.3 | 98.5 | 22 |
| CGD-CNN [34] | RGB-D | 97.8 | 96.6 | 19 |
| QQGNN [1] | RGB-D | 97.7 | 98.9 | 14.7 |
| HMT-Grasp (Ours) | RGB | 99.21 | 99.44 | 23.79 |
| D | 99.32 | 99.55 | 22.25 | |
| RGB-D | 99.55 | 99.55 | 24.38 | |
We first evaluated our method on the Cornell dataset using 5-fold cross-validation. Table I presents the results for both image-wise (IW) and object-wise (OW) splits. Our proposed approach achieved a detection accuracy of 99.55%, outperforming other state-of-the-art methods in both splits. The superior performance of HMT-Grasp can be attributed to its ability to effectively capture both local and global features, enabling more precise grasp detection even in challenging scenarios. This result demonstrates the robustness and effectiveness of our approach across various input data configurations in the Cornell dataset.
IV-C2 Grasping Performance on the Jacquard Dataset
| Years | Method | Input | Accuracy (%) |
| Morrison [9] | GGCNN2 | RGBD | 88.3 |
| Kumra [10] | GR-ConvNet | RGB-D | 91.0 |
| Wang [13] | TF-Grasp | RGB-D | 86.7 |
| Ours | HMT-Grasp | D | 93.0 |
| HMT-Grasp | RGB | 90.4 | |
| HMT-Grasp | RGB-D | 93.3 |
Data splitting for the Jacquard dataset varies across methods. For example, GGCNN2 uses 95% of the dataset for training, while TF-Grasp and GR-ConvNet use 90%, with minor differences in test splits. Additionally, distinctions between image-wise and object-wise splits are often unclear. To ensure fair comparisons, we reran all Jacquard experiments following the same protocol as state-of-the-art methods. The training and evaluation protocol is detailed in the Jacquard Dataset and Training section. As shown in Table II, our HMT-Grasp approach, using RGB-D input, achieved 93.3% accuracy, surpassing existing methods. The consistent improvement across different data splits reinforces the effectiveness of our hybrid architecture in capturing the critical features required for successful grasp prediction.
IV-C3 Grasping Performance on the OCID-Grasp Dataset
| Years | Method | Input | Accuracy (%) |
| Morrison [9] | GGCNN2 | D | 54.51 |
| GGCNN2 | RGB | 62.96 | |
| GGCNN2 | RGBD | 64.66 | |
| Kumra [10] | GR-ConvNet | D | 64.83 |
| GR-ConvNet | RGB | 68.12 | |
| GR-ConvNet | RGB-D | 68.69 | |
| Wang [13] | TF-Grasp | D | 41.63 |
| TF-Grasp | RGB | 66.20 | |
| TF-Grasp | RGB-D | 66.02 | |
| Ours | HMT-Grasp | D | 68.86 |
| HMT-Grasp | RGB | 70.16 | |
| HMT-Grasp | RGB-D | 70.62 |
To validate our method in complex environments, we evaluated it on the OCID-Grasp dataset. The results, presented in Table III, follow the settings outlined in the OCID Dataset and Training section. Our approach, using RGB-D input, outperformed other methods, demonstrating its superior ability to handle complex grasping scenarios, and showing greater adaptability compared to ViTs and CNNs. The superior performance in this cluttered environment underscores the enhanced feature extraction capability of the Mamba blocks, which balance local and global information to enable more effective grasping decisions in difficult scenarios.
IV-D Ablation Studies
To further investigate the impact of skip-connections and the Mamba block on grasp pose learning, we trained our network using 5-fold cross-validation on the Cornell dataset with image-wise splitting.
IV-D1 Ablation study for skip-connections
We firstly conducted experiments with and without skip-connections using our method across RGB, Depth, and RGB-D inputs. The detailed experimental results are presented in Table IV. The results show that incorporating skip-connections improves performance compared to omitting them, except for the Depth input.
| With Skip-connections | Without Skip-connections | |
| Depth(%) | 99.32 | 99.55 |
| RGB(%) | 99.21 | 99.20 |
| RGB+Depth(%) | 99.55 | 99.32 |
IV-D2 Ablation study for Mamba blocks
Following the same procedure as the skip-connection ablation study, we evaluated the effect of the Mamba block on the Cornell dataset. The results, shown in Table V, demonstrate that the inclusion of Mamba blocks leads to superior performance across Depth, RGB, and RGB-D inputs compared to the model without Mamba blocks.
| With Mamba Blocks | Without Mamba Blocks | |
| Depth(%) | 99.32 | 97.51 |
| RGB(%) | 99.21 | 98.53 |
| RGB+Depth(%) | 99.55 | 98.87 |
IV-E Simulation and Real Robot Experiments
In this section, we first conducted robot grasping experiments in the Gazebo using models.First, we performed 10 isolated experiments for each object by randomly spawning it within the robot’s workspace. The results, presented in Table VI, demonstrate the effectiveness of our method in grasping various objects. The reasons for failures can be attributed to the complexity of object shapes and material properties, such as sponges being soft and deformable, making it difficult for the grasp detection model to predict an optimal grasp configuration. As illustrated in Fig. 3, we also conducted 10 rounds of object grasping in cluttered scenarios for both simulation and real robots. In the case of simulation experiments, the robot was able to outperform the selected state-of-the-art with a 7% margin (65% vs 72%). The model’s ability to balance local and global feature extraction, thanks to the Mamba blocks, contributed significantly to its improved performance in complex scenarios. The real robot achieved a success rate comparable to the simulation results, with 67% successful grasps. The failures in real-world experiments were primarily due to object occlusion. In particular, many objects are partially occluded by others, leading to attempts to grasp objects at an inappropriate angle or position.
| Objects | Pear | lemon | sponge | can_coke | Peach |
| GR-ConvNet [10] | 80 | 75 | 85 | 80 | 75 |
| HMT-Grasp | 100 | 95 | 75 | 95 | 80 |
| Objects | mug | scissor | tea | banana | flash |
| GR-ConvNet [10] | 80 | 70 | 70 | 85 | 75 |
| HMT-Grasp | 70 | 95 | 80 | 85 | 90 |
V Conclusion
This paper presents a hybrid Mamba-Transformer model for robotic grasping in complex scenarios. By combining Mamba, Transformers, and CNNs, our method overcomes the limitations of ViTs and CNNs in visual grasp detection. We conducted extensive evaluations on the Cornell, Jacquard, and OCID-Grasp datasets, where our model consistently outperformed state-of-the-art methods, achieving an accuracy of 99.5% on the Cornell dataset. These results demonstrate the superior adaptability and precision of our approach. Additionally, real-world robotic experiments further validated the method’s effectiveness in dynamic and complex environments, highlighting its potential for practical applications. Future work could focus on incorporating real-time feedback mechanisms to dynamically adjust grasp plans during execution.

### v2

@@
 Robot grasping, whether handling isolated objects, cluttered items, or stacked objects, plays a critical role in industrial and service applications.
-However, current visual grasp detection methods based on Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) struggle to adapt across various grasping scenarios due to the imbalance between local and global feature extraction.
+However, current visual grasp detection methods based on Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) often struggle to adapt to diverse scenarios, as they tend to emphasize either local or global features exclusively, neglecting complementary cues.
 In this paper, we propose a novel hybrid Mamba-Transformer approach to address these challenges.
@@
 I Introduction
-Robotic grasping has become increasingly important in industrial production and service applications [1, 2].
-Tasks such as automated warehouse picking and household robots rely on accurate grasping to handle objects effectively.
-While early methods using hand-crafted features performed well in simple, controlled environments, modern robotic systems face more complex scenarios with diverse object shapes and unpredictable scenarios, making grasping a greater challenge [3].
-Deep learning techniques, particularly Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have greatly advanced computer vision tasks such as image classification, object detection, and segmentation [4, 5].
-Researchers have leveraged these techniques to improve robotic grasp detection by utilizing their ability to extract crucial visual information.
-Early CNN-based grasp detectors [6, 7, 8, 9, 10] demonstrated notable improvements over hand-crafted methods, adapting to a wider variety of objects.
-However, CNNs focus mainly on local feature extraction, often missing the global context needed for grasping in more complex scenarios.
-To address this issue, Vision Transformers [11, 12] have been introduced in grasping tasks, offering enhanced global feature representation by capturing information across the entire input image.
-For instance, Wang et al.
-[13] and Dong et al.
-[14] employed transformers to encode target scenes for robot grasping detection.
-While transformer-based approaches enhance global feature representation, they tend to decrease concern on local details as the model depth increases [15].
-A more recent approach, Vision Mamba (ViM), balances global and local information more effectively than ViTs by incorporating Visual State Space (VSS) models [16].
-While ViM improves spatial detail capture and focuses more on local features than transformers, its local feature extraction remains less refined than CNNs.
-In this work, we propose a hybrid Mamba-Transformer approach, integrating the strengths of Vision Mamba [17], transformers [18], and CNNs to enhance robotic grasp detection across various scenarios.
-By combining outputs from parallel CNNs and transformers as input to the Mamba encoder, our approach balances global and local feature representations, addressing the limitations of each model.
+Robotic grasping is a fundamental capability for autonomous manipulation in industrial automation, service robotics, and domestic applications [1, 2].
+Robots operate in real-world situations where objects vary in shape, size, and material properties and are randomly arranged and partially occluded.
+In such challenging conditions, robust and adaptive grasping strategies are needed to identify feasible grasp configurations despite significant visual and physical clutter [3].
+Conventional approaches to robotic grasping relied on analytical methods and hand-crafted features, which performed adequately under controlled conditions but failed to generalize to complex, unstructured environments.
+With the advent of deep learning, convolutional neural networks (CNNs) have become the predominant paradigm for extracting discriminative local features that are essential for grasp detection [4, 5, 6, 7, 8].
+However, in cluttered scenes, local features alone are often insufficient because grasping decisions depend critically on global contextual cues—such as spatial relationships among objects and prevalent occlusion patterns—which traditional CNNs may not adequately capture.
+Recently, transformer architectures have underscored their ability to model long-range dependencies and provide a more comprehensive contextual representation [9, 10].
+Nonetheless, these methods may inadvertently diminish the local spatial details—such as edges, corners, and shapes—that are crucial for accurate grasp detection [11].
+In contrast, recent advances in Mamba architectures extract both local and global features in a more balanced manner than ViTs and CNNs [12, 13], which tend to favor one over the other, thereby mitigating their limitations.
+To improve robot grasp detection performance in multiple scenarios, we propose a novel Mamba-Transformer architecture that leverages a state space model (SSM) to fuse and refine features extracted by parallel CNN and Transformer modules.
+By effectively integrating local detail with global contextual cues, our approach facilitates robust grasping detection across various scenes.
 Experimental results show that our method outperforms state-of-the-art techniques on multiple grasping datasets and delivers superior performance in both simulation and real-world robotic applications.
-The main contributions of this paper can be summarized as follows:
+The main contributions of this paper are summarized as follows:
 1.
-We propose a novel hybrid Mamba-Transformer approach that integrates the strengths of Vision Mamba, transformers, and CNNs to improve robotic grasp detection.
-To the best of our knowledge, we are the first to explore the performance of the Mamba model in robotic grasping tasks.
+We propose a hybrid Mamba-Transformer architecture that integrates parallel CNN and Transformer streams, leveraging their respective strengths to enable the simultaneous extraction of fine-grained local features and global contextual cues for grasp detection.
 2.
-We provide a comprehensive evaluation of our method on a diverse set of grasping datasets, demonstrating its effectiveness across various scenarios.
+We then present the Mamba encoder to further fuse local and global features, which refines fine local details while effectively integrating broader contextual information.
+3.
+We conduct comprehensive evaluations and rigorous ablation studies across multiple grasping datasets, achieving excellent performance.
+Additionally, we carry out extensive simulation and real robot experiments.
+The experimental results clearly demonstrate that our approach exhibits robust performance.
 II Related Work
@@
 III Method
-In this section, we introduce a hybrid Mamba-Transformer architecture for robotic grasp detection. The overall structure of the model is shown in Fig. 1.
-III-A Grasp Representation
-We define the grasp representation , based on prior works [13, 9], as key parameters describing a successful grasp.
-Using a parallel-plate gripper constrained to the - plane, the grasp is represented by:
-where are the center coordinates of the grasp rectangle, and are the width and height, is the rotation angle. For an image with channels and dimensions , the predicted grasp is:
-here, are the coordinates, is the rotation angle, is the grasp width, and is the grasp quality. To convert image coordinates into the robot’s reference frame, we apply a series of transformations:
-where transforms from image space to the camera frame using intrinsic parameters, and converts to the robot frame. The grasp information in the image is stored as a grasp map:
-where represent the angle, width, and quality for each pixel.
-The hybrid Mamba-Transformer framework analyzes the n-channel image to determine the optimal grasping pose, focusing on predicting the grasp location, orientation, and width.
-III-B Grasp Encoder
-The input tensor to the architecture has dimensions , where is the batch size, is the number of channels, and represent the height and width of the image.
-The tensor dimensions change as it passes through different layers of the model.
-III-B1 Stem-Downsample Block
-Initially, the input tensor is processed by the Stem-Downsample Block, consisting of two Conv2d layers, each followed by BatchNorm (BN) and SiLU activations:
-III-B2 Parallel Convolutional-Transformer Blocks
-Following the StemDown Block, the output is processed through parallel Transformer and Convolutional Blocks.
-Transformer Blocks: To capture global information, we employ the Swin Transformer block [11], which uses Window-based Multi-Head Self-Attention (W-MSA) and Shifted Window-based MSA (SW-MSA), followed by LayerNorm (LN) and Multi-Layer Perceptron (MLP).
-Convolutional Blocks: To refine local features, a DoubleConv block is used, consisting of Conv2D, BN, and ReSiLU activations, where ReSiLU combines the outputs of ReLU and SiLU activations:
-III-B3 Downsampling
-To merge the information from the Transformer and Convolutional Blocks, we first compute the average of and , denoted as .
-This is followed by a downsampling network consisting of MaxPooling and DoubleConv layers for further refinement:
-The resulting fused tensor undergoes an additional DoubleConv operation, ensuring efficient downsampling and feature extraction.
-III-B4 Mamba Blocks
-The fused and downsampled tensor with dimension is processed by the Mamba Block, inspired by the VM-UNet architecture [28].
-The Mamba Block consists of LayerNorm (LN), Linear Projection (LP), Conv2d, SiLU activation, and 2D-Selective-Scan (SS2D).
-The computation is expressed as:
-In this formulation, the input undergoes Layer Normalization before being processed by a series of operations, including a linear projection, convolution, and the SiLU activation.
-The resulting feature map is further refined using the 2D-Selective-Scan (SS2D) mechanism, followed by an element-wise combination of the two branches.
-A final residual connection is applied to retain the input information and improve gradient flow during training.
-This structure aims to effectively capture spatial details and enhance the overall feature representation for subsequent stages of the network.
-III-C Grasp Decoder
-After processing by the grasp encoder, the feature representation is passed to a DoubleConv block for further feature extraction.
-This DoubleConv block consists of Conv2D, Batch Normalization (BN), and ReSiLU activation layers, similar to the ones used in the grasp encoder.
-During this step, the input and output dimensions remain unchanged.
-The output of the DoubleConv block is combined with the output from the Mamba block.
-Mathematically, this can be expressed as:
-The first fused output () is passed through an upsampling block. After upsampling, the output with dimensions is processed by another DoubleConv block for refinement:
-The first upsampled output is then summed with the output from the parallel convolutional-transformer block in the grasp encoder, followed by an additional upsampling operation.
-This process ensures the preservation and refinement of spatial details through feature refinement, skip connections, and upsampling.
-The resulting tensor with dimensions is further upsampled to generate pixel-level heatmap images for grasp prediction.
-III-D Grasp Prediction
-In the final output of the HMT-Grasp detector, pixel-level heatmap images are generated for grasp representation.
-These include a grasp confidence map , a gripper angle map , and a gripper width map , all matching the input image size (224x224).
-Grasp posture estimation is formulated as a regression task, where the model learns the mapping to minimize the difference between predicted grasp heatmaps and ground truth , with as the input data.
-The loss function is defined as:
-where represents the number of samples, and denotes the ground truth for each sample.
-The optimal grasp location is determined by selecting the position with the highest grasp confidence , as defined by:
-where represents the chosen grasp location. The corresponding gripper angle and width are then extracted from the respective heatmaps.
+In this section, we present a hybrid Mamba-Transformer architecture for robotic grasp detection. The overall structure of the proposed model is illustrated in Fig. 1.
+III-A Grasp Encoder
+We first apply a Stem Block comprising two convolutional layers to downsample the input tensor.
+The resulting feature map is then fed into a parallel network architecture that integrates CNNs and Transformers, as shown in Fig.
+1.
+III-A1 Parallel Network
+To capture spatial relationships among objects, we employ the Swin Transformer block [24], which leverages window-based multi-head self-attention (WMSA) and shifted window-based multi-head self-attention (SWMSA).
+Although window-based attention extracted localized features, shifting windows promote cross-window interactions, thereby enhancing global modeling capability [11].
+However, Transformers alone tend to suppress fine-grained local features at increasing depths [11].
+To address this limitation, we incorporate convolutional neural networks (CNNs) within the Parallel Network.
+More specifically, we adopt a dual convolutional architecture [25] to capture localized details such as edges and shapes.
+This complementary design capitalizes on CNNs’ inherent translation invariance and inductive biases, leading to more robust representations for robotic grasp detection.
+III-A2 Mamba Blocks
+The Mamba Block employs a recursive state space model (SSM) that supports continuous information flow across the spatial domain.
+Two core mechanisms drive this design: implicit attention, which propagates long-range dependencies, and dynamic filtering, which adaptively fuses local and global features.
+When input features combine CNN-derived local details and transformer-based global context, Mamba’s dynamic filtering effectively balances these distinct feature types.
+Consequently, Vision Mamba achieves an optimal trade-off between preserving local detail and modeling broader structures.
+This capability enhances the overall feature representation capacity within the Parallel Network.
+Following the parallel feature extraction, we concatenate the outputs into of dimension .
+Inspired by the VM-UNet architecture [23], we feed this concatenated feature map into the mamba block, which comprises LayerNorm (LN), Linear Projection (LP), Conv2d, SiLU, and the 2D-Selective-Scan (SS2D) mechanism:
+where indicates element-wise multiplication. The residual connection ensures stable gradient flow, while SS2D promotes richer spatial feature extraction.
+III-B Grasp Decoder
+In the decoder phase, skip connections establish feature reuse by combining encoder outputs with corresponding decoder stages, as shown in Fig.
+1.
+We apply average pooling to the encoder feature maps, then upsample them by a factor of two.
+A subsequent DoubleConv operation reduces channel dimensionality:
+Through iterative refinement, the decoder progressively restores spatial resolution.
+Once feature maps reach , they are upsampled again to yield pixel-level heatmaps matching the original image resolution of .
+These heatmaps include a grasp confidence map , a gripper angle map , and a gripper width map .
+III-C Grasp Synthesis
+A robotic antipodal grasp can be defined by the tuple, where denotes the grasp center in pixel coordinates, is the in-plane rotation, is the gripper width, and indicates the predicted grasp success confidence.
+Once the pixel maximizing is found, its location is converted to Cartesian coordinates via known camera intrinsics and a calibrated transformation from the camera to the robot reference frame.
+Next, the depth information at is extracted to determine the object’s position along the optical axis.
+To ensure robust execution of the grasp, the system may also consider a small neighborhood around the selected pixel to accommodate variations in sensor noise or object geometry.
+Finally, the manipulator is instructed to close the gripper at the computed position and orientation, thereby completing the grasp action.
 IV Experiments
-In this section, we present a comprehensive evaluation of our method through experiments on the Cornell, Jacquard, and OCID-Grasp datasets.
-We assess the impact of skip connections and Mamba blocks on performance and compare our approach to state-of-the-art methods.
+In this section, we present a comprehensive evaluation of our method through experiments on the Cornellt [14], Jacquard [15], and OCID-Grasp [26] datasets.
+We evaluate the impact of skip connections, perform an ablation study on multiple components, and compare our approach against state-of-the-art methods.
 Additionally, we validate our method in both simulation and real-world robotic experiments.
@@
 We utilize 5-fold cross-validation, as in the Cornell dataset training, to train the networks and report average performance across the folds.
-The Adam optimizer is employed with a learning rate of 0.0004 and a batch size of 8.
-The network is trained for 70 epochs to achieve the final model weights.
+The Adam optimizer is employed with a learning rate of 0.001 and a batch size of 8.
+The network is trained for 80 epochs to achieve the final model weights.
 IV-B Evaluation Metrics
@@
 Angle Difference: The orientation difference between the predicted grasp and the ground-truth is less than 30.
-Jaccard Index: The Jaccard index between the predicted grasp and the ground-truth is greater than 25%, calculated as:
-here, and denote the predicted and ground-truth grasp rectangles, respectively, with the intersection and union represented as and .
+Jaccard Index: The Jaccard index between the predicted grasp and the ground-truth is greater than 25%.
 IV-C Performance Comparison with State-of-the-Art Methods
+IV-C1 Grasping Performance on the Cornell Dataset
 To demonstrate the effectiveness of our approach, we compared it against state-of-the-art methods under consistent training and testing settings across the Cornell, Jacquard, and OCID-Grasp datasets.
-IV-C1 Grasping Performance on the Cornell Dataset
+We first evaluated our method on the Cornell dataset using 5-fold cross-validation.
+Table I presents the results for both image-wise (IW) and object-wise (OW) splits.
+Our proposed approach achieved a detection accuracy of 99.55%, outperforming other state-of-the-art methods in both splits.
+The superior performance of HMT-Grasp can be attributed to its ability to effectively capture both local and global features, enabling more precise grasp detection even in challenging scenarios.
+This result illustrates the robustness and effectiveness of our approach across various input data configurations in the Cornell dataset.
+Additionally, the experimental result also demonstrates that our method achieves good real-time performance on the tested experimental equipment.
 | Method | Input | Accuracy (%) | Time (ms) | |
@@
 | QQGNN [1] | RGB-D | 97.7 | 98.9 | 14.7 |
-| HMT-Grasp (Ours) | RGB | 99.21 | 99.44 | 23.79 |
-| D | 99.32 | 99.55 | 22.25 | |
-| RGB-D | 99.55 | 99.55 | 24.38 | |
-We first evaluated our method on the Cornell dataset using 5-fold cross-validation.
-Table I presents the results for both image-wise (IW) and object-wise (OW) splits.
-Our proposed approach achieved a detection accuracy of 99.55%, outperforming other state-of-the-art methods in both splits.
-The superior performance of HMT-Grasp can be attributed to its ability to effectively capture both local and global features, enabling more precise grasp detection even in challenging scenarios.
-This result demonstrates the robustness and effectiveness of our approach across various input data configurations in the Cornell dataset.
+| HMT-Grasp (Ours) | RGB | 99.21 | 99.44 | 23.8 |
+| D | 99.32 | 99.55 | 22.3 | |
+| RGB-D | 99.55 | 99.55 | 24.8 | |
 IV-C2 Grasping Performance on the Jacquard Dataset
-| Years | Method | Input | Accuracy (%) |
+| Reference | Method | Input | Accuracy (%) |
 | Morrison [7] | GGCNN2 | RGBD | 88.3 |
@@
 IV-C3 Grasping Performance on the OCID-Grasp Dataset
-| Years | Method | Input | Accuracy (%) |
-| Morrison [9] | GGCNN2 | D | 54.51 |
-| GGCNN2 | RGB | 62.96 | |
-| GGCNN2 | RGBD | 64.66 | |
-| Kumra [10] | GR-ConvNet | D | 64.83 |
-| GR-ConvNet | RGB | 68.12 | |
-| GR-ConvNet | RGB-D | 68.69 | |
-| Wang [13] | TF-Grasp | D | 41.63 |
-| TF-Grasp | RGB | 66.20 | |
-| TF-Grasp | RGB-D | 66.02 | |
-| Ours | HMT-Grasp | D | 68.86 |
-| HMT-Grasp | RGB | 70.16 | |
-| HMT-Grasp | RGB-D | 70.62 |
+| Reference | Method | Input | Accuracy (%) |
+| Morrison [7] | GGCNN2 | D | 69.3 |
+| GGCNN2 | RGB | 65.7 | |
+| GGCNN2 | RGBD | 67.6 | |
+| Kumra [8] | GR-ConvNet | D | 70.0 |
+| GR-ConvNet | RGB | 68.5 | |
+| GR-ConvNet | RGB-D | 72.3 | |
+| Wang [9] | TF-Grasp | D | 72.0 |
+| TF-Grasp | RGB | 58.9 | |
+| TF-Grasp | RGB-D | 61.8 | |
+| Ours | HMT-Grasp | D | 73.5 |
+| HMT-Grasp | RGB | 76.5 | |
+| HMT-Grasp | RGB-D | 74.1 |
 To validate our method in complex environments, we evaluated it on the OCID-Grasp dataset.
 The results, presented in Table III, follow the settings outlined in the OCID Dataset and Training section.
-Our approach, using RGB-D input, outperformed other methods, demonstrating its superior ability to handle complex grasping scenarios, and showing greater adaptability compared to ViTs and CNNs.
+Our approach, using RGB input, outperformed other methods, demonstrating its superior ability to handle complex grasping scenarios, and showing greater adaptability compared to ViTs and CNNs.
+As shown in Fig 2, the grasping detection results demonstrates our proposed grasp detection method can achieve a good performance across multiple objects scenarios.
 The superior performance in this cluttered environment underscores the enhanced feature extraction capability of the Mamba blocks, which balance local and global information to enable more effective grasping decisions in difficult scenarios.
@@
 The detailed experimental results are presented in Table IV.
-The results show that incorporating skip-connections improves performance compared to omitting them, except for the Depth input.
+The study results indicate that incorporating skip connections improves grasp detection performance, particularly for RGB-based inputs, achieving a 2.3 improvement compared to the absence of skip connections.
 | With Skip-connections | Without Skip-connections | |
-| Depth(%) | 99.32 | 99.55 |
-| RGB(%) | 99.21 | 99.20 |
-| RGB+Depth(%) | 99.55 | 99.32 |
-IV-D2 Ablation study for Mamba blocks
-Following the same procedure as the skip-connection ablation study, we evaluated the effect of the Mamba block on the Cornell dataset.
-The results, shown in Table V, demonstrate that the inclusion of Mamba blocks leads to superior performance across Depth, RGB, and RGB-D inputs compared to the model without Mamba blocks.
-| With Mamba Blocks | Without Mamba Blocks | |
-| Depth(%) | 99.32 | 97.51 |
-| RGB(%) | 99.21 | 98.53 |
-| RGB+Depth(%) | 99.55 | 98.87 |
+| Depth(%) | 73.5 | 73.1 |
+| RGB(%) | 76.5 | 74.2 |
+| RGB-D(%) | 74.1 | 73.4 |
+IV-D2 Ablation study for different components
+To evaluate the contributions of the CNN, Transformer, and Mamba modules in our proposed method, we conduct an ablation study on the OCID-Grasp dataset.
+The results are presented in Table V.
+As shown in the first row, using CNN-only achieves accuracy of 73.9%.
+When replacing the CNN with only the Transformer and Mamba modules, the accuracy drops to 71.2%, indicating that the convolutional network captures essential local grasping features, such as the textures and edge, as shown in Fig.
+3 that significantly contribute to detection performance.
+When combining the CNN with the Mamba module, the accuracy remains at 73.9%, showing no substantial improvement compared to using only the CNN or Mamba.
+As shown in Fig.
+3, the quality heatmaps of Mamba-only and CNN-only models indicate that both focus more on local details compared to the Transformer-only model.
+This also suggests that while Mamba effectively enhances sequence modeling, it lacks the capability to fully compensate for the absence of strong global representations, which are crucial in complex, multi-object grasping scenarios where spatial relationships between objects and background diversity play a significant role.
+In contrast, the combination of the CNN and Transformer results in a slight performance drop compared to the standalone components.
+This outcome underscores the importance of the Mamba module in balancing local and global representations.
+The Mamba module refines feature extraction by integrating local information from the CNN and global spatial awareness from the Transformer, ensuring that the model does not overly prioritize one type of representation at the expense of the other.
+Finally, our method, which integrates CNN, Transformer, and Mamba, achieves the highest accuracy of 76.5%, demonstrating Mamba’s effectiveness in furthering refine local and global features for enhanced grasp detection.
+These results validate our architecture.
+CNN primarily captures fine local details, the Transformer focuses on extracting global context, and Mamba effectively fuses and enhances both, ultimately improving overall performance.
+| Ablation Settings | Accuracy(%) | ||
+| CNN | Transformer | Mamba | |
+| ✓ | 73.9 | ||
+| ✓ | 74.0 | ||
+| ✓ | 74.1 | ||
+| ✓ | ✓ | 72.7 | |
+| ✓ | ✓ | 71.2 | |
+| ✓ | ✓ | 73.9 | |
+| ✓ | ✓ | ✓ | 76.5 ( Ours) |
 IV-E Simulation and Real Robot Experiments
-In this section, we first conducted robot grasping experiments in the Gazebo using models.First, we performed 10 isolated experiments for each object by randomly spawning it within the robot’s workspace.
+In this section, we first conducted robot grasping experiments in the Gazebo using models, as shown in Fig.
+5.
+First, we performed 20 isolated experiments for each object by randomly spawning it within the robot’s workspace.
 The results, presented in Table VI, demonstrate the effectiveness of our method in grasping various objects.
-The reasons for failures can be attributed to the complexity of object shapes and material properties, such as sponges being soft and deformable, making it difficult for the grasp detection model to predict an optimal grasp configuration.
-As illustrated in Fig.
-3, we also conducted 10 rounds of object grasping in cluttered scenarios for both simulation and real robots.
-In the case of simulation experiments, the robot was able to outperform the selected state-of-the-art with a 7% margin (65% vs 72%).
-The model’s ability to balance local and global feature extraction, thanks to the Mamba blocks, contributed significantly to its improved performance in complex scenarios.
-The real robot achieved a success rate comparable to the simulation results, with 67% successful grasps.
-The failures in real-world experiments were primarily due to object occlusion.
-In particular, many objects are partially occluded by others, leading to attempts to grasp objects at an inappropriate angle or position.
+The results show that HMT-Grasp generally outperforms other methods across most objects.
+It achieved 100% success on the pear, surpassing GR-ConvNet (80%) and TF-Grasp (85%).
+It also performed well on can_coke (95%), exceeding all other methods.
+Although HMT-Grasp generally performed better, there were cases where the improvement was marginal.
+For example, HMT-Grasp and GR-ConvNet both reached 85% on the banana, and the performance difference for the peach was only .
+The reasons for failures can be attributed to the complexity of object shapes, material and collision mesh properties.
+For example, objects like the mug had lower success rates for both methods.
+The challenge with grasping this object lies in the difficulty of detecting stable grasp points, as small surface areas reduce friction and stability.
+We conducted real-robot experiments in lightly, moderately, and heavily cluttered environments with complex backgrounds, as illustrated in Fig.
+4.
+In the lightly cluttered setting, we tested twenty groups of five randomized objects per method to ensure diverse grasping scenarios.
+In the moderately cluttered setting, we conducted ten groups of ten randomized objects per method to evaluate grasping performance across various backgrounds.
+In the highly cluttered setting, we ran five experimental groups, each aiming to include twenty distinct objects to assess robustness under dense clutter.
+We compute the success rate as the ratio of successful grasps to the total number of grasp attempts.
+The results are presented in Table VII.
+Overall, our approach consistently showed better performance compared to the baseline methods.
+However, accuracy decreased from lightly to heavily cluttered scenarios for two primary reasons: (1) insufficient friction between the gripper jaws and certain smooth-surfaced objects, resulting in grasp failures; and (2) the inaccurate grasp prediction led to contacting the gripper with the neighboring objects, resulting in a grasp failure.
 | Objects | Pear | lemon | sponge | can_coke | Peach |
 | GR-ConvNet [8] | 80 | 75 | 85 | 80 | 75 |
+| TF-Grasp [9] | 85 | 70 | 70 | 85 | 70 |
 | HMT-Grasp | 100 | 95 | 75 | 95 | 80 |
-| Objects | mug | scissor | tea | banana | flash |
+| Objects | mug | scissor | tea_box | banana | flash |
 | GR-ConvNet [8] | 80 | 70 | 70 | 85 | 75 |
+| TF-Grasp [9] | 75 | 85 | 80 | 75 | 80 |
 | HMT-Grasp | 70 | 95 | 80 | 85 | 90 |
+| Models | Success Rate () | ||
+| Light Clutter | Moderate Clutter | High Clutter | |
+| (5 Objects) | (10 Objects) | (20 Objects) | |
+| GR-ConvNet [8] | 80 | 74 | 72 |
+| TF-Grasp [9] | 82 | 75 | 70 |
+| HMT-Grasp | 88 | 82 | 79 |
 V Conclusion
 This paper presents a hybrid Mamba-Transformer model for robotic grasping in complex scenarios.
-By combining Mamba, Transformers, and CNNs, our method overcomes the limitations of ViTs and CNNs in visual grasp detection.
+By combining Mamba, Transformers, and CNNs, our method improves the performance of ViTs and CNNs in visual grasp detection.
 We conducted extensive evaluations on the Cornell, Jacquard, and OCID-Grasp datasets, where our model consistently outperformed state-of-the-art methods, achieving an accuracy of 99.5% on the Cornell dataset.
 These results demonstrate the superior adaptability and precision of our approach.
-Additionally, real-world robotic experiments further validated the method’s effectiveness in dynamic and complex environments, highlighting its potential for practical applications.
-Future work could focus on incorporating real-time feedback mechanisms to dynamically adjust grasp plans during execution.
+Additionally, real-world robotic experiments further validated the method’s effectiveness in complex environments, highlighting its potential for practical applications.
+Future work could focus on incorporating real-time feedback mechanisms to adjust grasp plans dynamically during execution.